OpenStudy (anonymous):

Which of the following is a polynomial with roots 4, –2, and 2i

OpenStudy (anonymous):

the one that looks like \((x-4)(x+2)(x+2i)(x-2i)\)

OpenStudy (anonymous):

that is , the one that looks like that stuff after you multiply out

OpenStudy (anonymous):

these are the possible answers x3 – 4x2 + 4x – 16 x3 + 4x2 + 4x + 16 x3 + 2x2 + 4x + 8 x3 – 2x2 + 4x – 8

OpenStudy (anonymous):

(x + 4)(x - 2i)(x + 2i) = (x + 4)(x² + 4) = x³ + 4x² + 4x + 16

OpenStudy (anonymous):

see here same question is here http://answers.yahoo.com/question/index?qid=20110424143827AA7IXkt

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

@best.shakir its 4, -2, and 2i not -2i

OpenStudy (anonymous):

what about -2 as a zero??

OpenStudy (anonymous):

must be a typo in the question solution should be a polynomial of degree 4, not 3

OpenStudy (anonymous):

so do u think the last one x3 – 2x2 + 4x – 8 wud be the answer

OpenStudy (anonymous):

never mind i have the answer